OpenStudy (anonymous):

Solve each equation on the interval [0, 2pi]. 4sin^2-3=0 rewrite this equation so that it is expressed in terms of sin^2x=a, where a is a nonzero number.

OpenStudy (jdoe0001):

\(\bf 4sin^2(x)-3=0 \ \ ?\)

OpenStudy (anonymous):

\[\sin \Theta = \pm \frac{ \sqrt{3} }{ 2}\]

OpenStudy (anonymous):

Just solve for all solutions on the unit circle regarding the angle \[\pi / 3\]

OpenStudy (anonymous):

what do you mean by that? I really don't understand how to do this problem.

OpenStudy (anonymous):

Here is a unit circle.. find out the rest of the angles